OpenStudy (anonymous):

find the slope of these points (9,3) and (-3-2)

OpenStudy (anonymous):

.4166666666666667

OpenStudy (anonymous):

is that a positive or negative 2

OpenStudy (anonymous):

????

OpenStudy (anonymous):

negative -2

OpenStudy (anonymous):

the answer is .4167

OpenStudy (anonymous):

-5/-12

OpenStudy (anonymous):

can't have slope in decimals

OpenStudy (anonymous):

u need the x and y

OpenStudy (anonymous):

lol your right its been a long time since ive done slopes do not go off my answer, loco85 knows what he/she is talking about

OpenStudy (anonymous):

(-2-3)/(-3-9)=(-5/-12)

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

can you help me to find the slope of these points 3,8 and 3,-3

OpenStudy (anonymous):

do u know the slope formula?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no no se

OpenStudy (anonymous):

\[m=y2-y1/x1-x2\]

OpenStudy (anonymous):

(-3-8)/(3-3)=-11/0 which doesn't make sense so there is no slope

OpenStudy (anonymous):

ok and can you help me to use these equation to find the slope and y intercept of -3x+5y=10